two wires A and B of the same material and equal lengths but of radii r and soldered coaxially the free end of B is twisted by an angle the ratio of the twist at the junction and angle phi is


1
Expert's answer
2021-09-19T18:31:39-0400

"\text{Angle of twist is inversely proportional to the fourth square of the radius}\\
\theta=\dfrac{k}{r^4}\\
\text{k is constant}\\
\theta_A=\dfrac{k}{r_A^4}\\
\theta_B=\dfrac{k}{r_B^4}\\
\dfrac{\theta_A}{\theta_B}=(\dfrac{r_B}{r_A})^4"